# Split & Merge: Unlocking the Potential of Visual Adapters via Sparse Training

### 05 Dec 2023 v1

@@ 3.3 Stochastic activation tuning
Figure 3: Hierarchical Mixture of Sparse Adapters. We preserve the down-projection layer in the adapter as a dense matrix, while only applying sparse splitting to the up-projection layer.
With multiple sparse adapters, each can be treated as an expert, allowing for mixed training. Traditional MoE methods [61, 78] involve a routing mechanism, introducing additional computation and load imbalance. Considering that Adapter Tuning itself serves as a parameter-efficient method, and inspired by  [80, 68], in the training process of MoSA, we also adopt a completely stochastic activation mechanism, which has been demonstrated to be simple yet effective in the following experiments. Given two partitioned parameter sets $\mathscr{W}^{\text{down}}$ and $\mathscr{W}^{\text{up}}$, for each batch of data, we randomly sample one module $\mathcal{W}_{i}^{\text{down}}\in\mathscr{W}^{\text{down}},\mathcal{W}_{j}^{\text{up}}\in\mathscr{W}^{\text{up}}$ from each set to form an adapter $\mathcal{W}=\{\mathcal{W}_{i}^{\text{down}},\mathcal{W}_{j}^{\text{up}}\}$, as shown in Eq. 2. This activation mechanism ensures consistency in parameters and computational load with standard Adapter Tuning during the training process. Additionally, activating different sparse modules each time enables the model to learn different representations, thereby increasing the model capacity.
Hierarchical sparse strategy. Although the MoE system enhances the model capacity, the sparse activation also reduces the amount of data seen by individual experts, resulting in suboptimal performance, especially when training data for downstream tasks is limited. The pre-pruning and sparse training methods can mitigate this issue to a certain extent, and building upon this, we further propose a hierarchical sparse strategy, as illustrated in Fig. 3. The adapter module consists of two parts: the down-projection layer and the up-projection layer. We keep the down-projection layer as a dense matrix, i.e. $\mathscr{W}^{\text{down}}=\mathcal{W}^{\text{down}}$, and only introduce sparsity in the up-projection layer, i.e. $\mathscr{W}^{\text{up}}=\{\mathcal{W}_{i}^{\text{up}}\}_{i=1}^{N}$. This adaptation process can be expressed as:
@@
$\begin{split}\mathcal{L}(x,y)=\text{CE}(p_{1},y)+\frac{\alpha}{2}\left(\text{KL}(p_{1}\|p_{2})\right.+\\ \left.\text{KL}(p_{2}\|p_{1})\right)+\beta\sum_{i=1}^{L/2}\text{MSE}(f_{1}^{i},f_{2}^{i})\end{split}$ (6)
where $p_{1},p_{2}$ represent the predicted probabilities after two stochastic forward passes of $x$, and $f_{i}^{i},f_{2}^{i}$ represent the intermediate features in the $i$-th layer. Here, CE is the cross-entropy loss, KL is the Kullback–Leibler divergence and MSE is the mean square error. $\alpha$ and $\beta$ are hyper-parameters that control the weight of regularization terms.
@@ 3.4 Jigsaw-like adapter merging
During the inference stage, we merge the trained multiple sparse adapters like puzzle pieces into a complete adapter. The merging process can be expressed as follows:
@@
Here, $\mathcal{M}_{i}$ represents the sparse mask from Eq. 4. After the merge, the inference phase can be represented as:
@@ 4.1 Experimental setup
  • Linear probing: fix the model backbone and update only a single-layer classifier.
@@
  • AdaptFormer [6]: insert new bottleneck modules with residual connections inside Transformer blocks.
  • SparseAdapter [24]: prune the standard adapter before tuning and update the remaining parameters.
@@
Table 1: Results on FGVC with ViT-B/16 backbone pre-trained on ImageNet-21K.
Method Dataset VTAB GICD
Natural Specialized Structured Mean Acc. (%) Mean Params. (M) CIFAR-100 Aircraft Food-101 Mean Acc. (%) Mean Params. (M)
Full fine-tuning 75.88 83.36 47.64 68.96 85.80 89.12 70.93 90.96 83.67 85.80
Linear probing 68.93 77.16 26.84 57.64 0.00 85.95 45.06 88.14 73.05 0.00
BitFit [74] 73.30 78.25 44.09 65.21 0.10 91.69 68.71 89.59 83.33 0.10
AdaptFormer [6] 78.42 83.41 49.17 70.33 0.30 91.86 71.71 90.89 84.82 1.19
SparseAdapter [24] 77.58 81.99 48.26 69.28 0.08 91.20 67.15 89.37 82.57 0.30
MoSA (Ours) 79.86 84.03 50.28 71.39 0.30 92.22 72.14 91.17 85.18 1.19
Table 2: Results on VTAB and GICD with ViT-B/16 backbone pre-trained on ImageNet-21K.
4.2 Main Results
We provide a comprehensive evaluation of the effectiveness of our MoSA by comparing it with other baselines across 3 sets of up to 27 different datasets. We also combine five FGVC datasets together to form a complex multi-task benchmark, validating the enhancement of model capacity through the mixed training of multiple sparse experts. More experiments on different model architectures, adapter structures and PEFT methods (e.g., LoRA) can be found in Appendix. In the following experiments, Top-1 accuracy (%) is used to evaluate the performance of the methods on the respective datasets, and the number (M) of extra parameters (trainable parts excluding the classifier) is used to assess the efficiency of the methods. The best accuracy is highlighted in bold, while the second one is underlined.
@@ 4.2 Main Results
Fine-grained classification performance. We first evaluate the effectiveness of our method in transfer learning on 5 fine-grained visual classification tasks with ViT-B/16 [13] backbone. As shown in Tab. 1, our MoSA beats other fine-tuning baselines, including full fine-tuning, by a significant margin. Across the 5 downstream tasks, MoSA achieves an average accuracy of 89.90%, surpassing full fine-tuning and AdaptFormer [6] by 1.36% and 1.32%, respectively, while maintaining the same number of trainable parameters as [6]. Interestingly, SparseAdapter [24], with fewer trainable parameters just through pruning, outperforms the standard adapter by 0.12% on average, demonstrating the parameter redundancy in adapters for visual tasks. However, the high sparsity level results in ineffective utilization of most parameters in the adapter, limiting the overall performance gain. Our MoSA increases model capacity by mixed training of multiple sparse adapters, achieving an additional improvement of 1.20%. Experiments show that our method enhances the performance of existing methods without introducing extra parameters or computation, maximizing the potential of adapters.
@@
Low-resource visual adaptation performance. We also compare our method with other fine-tuning approaches on VTAB, which contains 19 diverse downstream tasks with only 1000 training samples per task, making it extremely challenging. Previous stochastically activated MoE methods, like  [80], suffer from severe performance degradation when training data is insufficient. However, our MoSA overcomes this issue with its hierarchical sparse strategy. The results in  Tab. 2 demonstrate that MoSA outperforms all other baselines by updating only 0.35% (0.30M in 85.80M) of the pre-trained backbone parameters. Specifically, across 3 domains in VTAB, MoSA surpasses full fine-tuning by 3.98%, 0.67% and 2.64%, while outperforming the second-best AdaptFormer by 1.44%, 0.62% and 1.11%, providing strong evidence for the effectiveness of our hierarchical sparse design.
@@
General large-scale classification performance. To further evaluate the generality of our method, we compare MoSA with other fine-tuning methods on 3 general classification tasks, as shown in Tab. 2. MoSA outperforms all baselines, including full fine-tuning, on all 3 datasets. Specifically, compared to AdaptFormer and full fine-tuning, MoSA achieves an average accuracy improvement of 1.51% and 0.36%, respectively, without introducing any additional parameters. It is worth noting that SparseAdapter performs poorly on this benchmark, exhibiting an accuracy drop of 2.25% compared to the standard adapter. This is attributed to the fact that the three datasets in GICD contain relatively sufficient training data, mitigating the advantages of sparsity. However, our method, through the combination of sparse training and MoE, demonstrates robust performance in scenarios with both abundant and limited training data, outperforming SparseAdapter by 2.61%, which proves the soundness of our design.
Method FGVC-multi (%) Params. (M)
Full fine-tuning 77.80 85.80
Linear probing 67.06 0.00
BitFit [74] 75.96 0.10
AdaptFormer [6] 78.28 2.42
SparseAdapter [24] 78.08 0.60
MoSA (Ours) 78.59 2.42
Table 3: Results on FGVC-multi with ViT-B/16 backbone pre-trained on ImageNet-21K.
Multi-task classification performance. In contrast to typical classification tasks, we mix the 5 datasets in FGVC to form a more challenging multi-task classification scenario, showcasing the superiority of our approach. Tab. 3 shows the comparison results against other baselines, with MoSA outperforming full fine-tuning by 0.79%, demonstrating the outstanding performance of sparse MoE design in multi-task settings. AdaptFormer and SparseAdapter achieve 78.28% and 78.08% accuracy respectively, while MoSA beats them by 0.31% and 0.51%, indicating that standalone sparse training takes a toll on performance, whereas its combination with MoE leads to improvements.
@@ 4.3 Ablation Study
We conduct comprehensive ablation studies to verify the effectiveness of each component in our MoSA design. All ablation experiments are performed on FGVC with ViT-B backbone, and the performances of different strategies are measured using mean accuracy over the datasets.
@@
Table 4: Effectiveness of hierarchical sparse strategy.
Hierarchical sparse strategy. The MoE system increases the model capacity, while the sparse gating mechanism also leads to a decrease in the amount of data each expert encounters, resulting in suboptimal performance particularly when training data for downstream tasks is limited. In order to address this issue, we propose a hierarchical sparse strategy, and the results in Tab. 4 demonstrate the effectiveness of this design. Applying the sparse strategy to both the down-projection and up-projection layer only achieves an accuracy of 88.91%, slightly (0.33%) ahead of Adaptformer (dense down-projection and up-projection layer). However, with the hierarchical sparse strategy, we preserve the down-projection layer as a dense matrix while sparsely splitting the up-projection layer. In this way, MoSA outperforms Adaptformer by a large margin (89.90% vs. 88.58%), demonstrating the importance of this hierarchical strategy.
@@
Table 5: Effectiveness of consistency regularization.
Consistency regularization. The mechanism of stochastic activation may lead to significant discrepancies among different experts, which can have adverse effects on the sparse module merging. In  Tab. 5, we investigate the impact of regularization constraints in the training of MoSA. Without any regularization, our method achieves an accuracy of 88.93%. Applying consistency regularization on the final outputs of the model results in an improvement of 0.32%. And deep alignment of features extracted by different experts can further improve the accuracy by 0.65%.
@@
Table 6: Different mechanisms of inference.
Merging vs. ensembling. During inference, we merge multiple sparse adapters to form a complete adapter. However, [80] has pointed out that stochastic activation during inference can also achieve good performance. Therefore, we compare various inference methods, including fixed activation, stochastic activation, logits ensembling, and parameter merging. Results are presented in  Tab. 6. It can be observed that fixed activation has the lowest accuracy, reaching only 88.41% on average. In comparison, stochastic activation achieves an improvement of 0.85%. Logits ensembling also leads to a certain improvement over fixed activation, but the increase of only 0.21% comes at the cost of $N$ times the computational complexity ($N$ refers to the number of experts), significantly reducing the inference speed. Finally, our parameter merging method achieves the best performance without any additional computation, further enhancing 0.63% over stochastic activation.
@@
Table 7: Impact of expert number.
Expert number. To evaluate the compatibility between sparse training and MoE, we vary the number of splits in the adapters of MoSA during training, ranging from 1 to 5 and 8. The results are presented in Tab. 7. When the number of splits is set to 1, our method degrades to AdaptFormer. We can see that when the number of experts is between 2 and 5, indicating a sparsity level between 20% and 50% for the adapters, MoSA consistently achieves good performance. However, when the number of experts increases to 8, the performance experiences a noticeable decline due to excessive sparsity.
Figure 4: Impact of bottleneck dimensions of the adapter.
Adapter bottleneck dimension. Fig. 4 shows the impact of bottleneck dimensions of the adapter in AdaptFormer and MoSA. Overall, the performance shows an increasing trend followed by a decline as the number of trainable parameters increases. Across all bottleneck dimensions, MoSA consistently outperforms AdaptFormer, and our method can even surpass full fine-tuning with a bottleneck dimension of only 16. It is worth noting that the performance of AdaptFormer starts to decline when the bottleneck dimension exceeds 64, while in our method, this turning point occurs at 128. This indicates that our design can more effectively utilize all parameters within the adapter.

### 23 Mar 2024 v2

@@ 3.3 Stochastic Activation Tuning
With multiple sparse adapters, each can be treated as an expert, allowing for mixed training. Traditional MoE methods [67, 84] involve a routing mechanism, introducing additional computation and load imbalance. Considering that Adapter Tuning itself serves as a parameter-efficient method, and inspired by  [86, 73], in the training process of MoSA, we also adopt a completely stochastic activation mechanism, which has been demonstrated to be simple yet effective in the following experiments. Given two partitioned parameter sets $\mathscr{W}^{\text{down}}$ and $\mathscr{W}^{\text{up}}$, for each batch of data, we randomly sample one module $\mathcal{W}_{i}^{\text{down}}\in\mathscr{W}^{\text{down}},\mathcal{W}_{j}^{\text{up}}\in\mathscr{W}^{\text{up}}$ from each set to form an adapter $\mathcal{W}=\{\mathcal{W}_{i}^{\text{down}},\mathcal{W}_{j}^{\text{up}}\}$, as shown in Equation 2. This activation mechanism ensures consistency in parameters and computational load with standard Adapter Tuning during training. Additionally, activating different sparse modules each time enables the model to learn different representations, thereby increasing the model capacity.
Hierarchical sparse strategy. Although the MoE system enhances the model capacity, the sparse activation also reduces the amount of data seen by individual experts, resulting in suboptimal performance, especially when training data for downstream tasks is limited. The pre-pruning and sparse training methods can mitigate this data dilution issue to a certain extent. Building upon this, we further propose a hierarchical sparse strategy. The adapter module consists of two parts: the down-projection layer and the up-projection layer. We keep the down-projection layer as a dense matrix, i.e. $\mathscr{W}^{\text{down}}=\mathcal{W}^{\text{down}}$, and only introduce sparsity in the up-projection layer, i.e. $\mathscr{W}^{\text{up}}=\{\mathcal{W}_{i}^{\text{up}}\}_{i=1}^{N}$. This adaptation process can be expressed as:
@@
$\mathcal{L}(x,y)=\text{CE}(p_{1},y)+\frac{\alpha}{2}\left(\text{KL}(p_{1}\|p_{2})+\text{KL}(p_{2}\|p_{1})\right)+\beta\sum_{i=1}^{L/2}\text{MSE}(f_{1}^{i},f_{2}^{i})$ (6)
where $p_{1},p_{2}$ represent the predicted probabilities after two stochastic forward passes of $x$, and $f_{i}^{i},f_{2}^{i}$ represent the intermediate features in the $i$-th layer. Here, CE is the cross-entropy loss, KL is the Kullback–Leibler divergence and MSE is the mean square error. $\alpha$ and $\beta$ are hyper-parameters that control the weight of regularization terms, which are simply set to $1.0$ in our main experiments.
@@ 3.4 Jigsaw-like Adapter Merging
During inference, we merge the trained multiple sparse adapters like puzzle pieces into a complete adapter, as illustrated in Figure 2. The merging process can be expressed as follows:
@@
Here, $\mathcal{M}_{i}$ represents the sparse mask for $\mathcal{W}_{i}$ from Equation 4, and $\overline{\mathcal{W}}$ represents the merged complete projection layer weight. After merging, the inference phase can be represented as:
@@ 4 Experiments
We evaluate MoSA across a series of downstream recognition tasks, spanning various model scales, architectures, and PEFT methods. We first describe our experimental setup in Section 4.1. Subsequently, we present the main experimental results in Section 4.2. Furthermore, we demonstrate the performance of MoSA applied to different backbone scales and PEFT methods in Section 4.3. In addition, we conduct extensive ablation experiments in Section 4.4 to verify the effectiveness of each component in our design. Finally, we provide t-SNE visualization results in Section 4.5.
@@ 4.1 Experimental Setup
  • Linear probing: fix the model backbone and only update the classifier.
@@
  • Visual prompt tuning (VPT) [36]: add learnable embeddings as prompts to modify the input, in two versions: shallow (insert prompts only into the first layer) and deep (introduce prompts at every layer).
  • AdaptFormer [7]: insert bottleneck modules with residual connections to the feed-forward network (FFN) of each Transformer layer.
  • SparseAdapter [30]: prune the standard adapter before tuning and update the remaining parameters via masks.
@@
4.2 Main Results
We provide a comprehensive evaluation of the effectiveness of our MoSA by comparing it with other baselines across 3 sets of up to 27 different datasets. In the following experiments, Top-1 accuracy (%) is used to evaluate the performance of the methods on the respective datasets, and the number (M) of extra parameters (trainable parts excluding the classifier) is used to assess the efficiency of the methods. The best accuracy is highlighted in bold, while the second one is underlined. The results of our method is highlighted with a red background.
Table 1: Results on FGVC with ViT-B/16 backbone pre-trained on ImageNet-21K
@@ 4.2 Main Results
Fine-grained classification performance. We first evaluate the effectiveness of our method on 5 widely used fine-grained visual classification tasks with ViT-B/16 [15] backbone. As shown in Table 1, our MoSA beats other baselines, including full fine-tuning, by a significant margin. Across the 5 downstream tasks, MoSA achieves an average accuracy of 89.90%, surpassing full fine-tuning and AdaptFormer [7] by 1.36% and 1.32%, respectively, while maintaining the same number of trainable parameters as [7]. Interestingly, SparseAdapter [30], with fewer trainable parameters just through pruning, outperforms the standard adapter by 0.12% on average, demonstrating the parameter redundancy in adapters for visual tasks. However, the high sparsity level results in ineffective utilization of most parameters in the adapter, limiting the overall performance gain. MoSA increases model capacity by mixed training of multiple sparse adapters, achieving an additional improvement of 1.20%. Experiments show that our method enhances the performance of existing methods without introducing extra parameters or computation, maximizing the potential of adapters.
Table 2: Results on VTAB with ViT-B/16 backbone pre-trained on ImageNet-21K
Method Dataset Natural Specialized Structured Mean Acc. (%) Mean Params. (M)
Full fine-tuning 75.88 83.36 47.64 68.96 85.80
Linear probing 68.93 77.16 26.84 57.64 0.00
BitFit [80] 73.30 78.25 44.09 65.21 0.10
VPT-shallow [36] 76.81 79.66 46.98 64.85 0.11
VPT-deep [36] 78.48 82.43 54.98 69.42 0.98
AdaptFormer [7] 78.42 83.41 49.17 70.33 0.30
SparseAdapter [30] 77.58 81.99 48.26 69.28 0.08
MoSA (Ours) 79.86 84.03 50.28 71.39 0.30
@@
Low-resource visual adaptation performance. We also compare our method with other fine-tuning approaches on VTAB, which contains 19 diverse downstream tasks with only 1000 training samples per task, making it extremely challenging. Previous stochastically activated MoE methods, like  [86], suffer from severe performance degradation when training data is insufficient. However, our MoSA overcomes this issue with its sparse training paradigm and hierarchical sparse strategy. The results in Table 2 demonstrate that MoSA outperforms all other baselines by updating only 0.35% (0.30M in 85.80M) of the pre-trained backbone parameters. Specifically, across 3 domains in VTAB, MoSA surpasses full fine-tuning by 3.98%, 0.67% and 2.64%, while outperforming the second-best AdaptFormer by 1.44%, 0.62% and 1.11%, providing strong evidence for the effectiveness of our design.
Table 3: Results on GICD with ViT-B/16 backbone pre-trained on ImageNet-21K
Method Dataset CIFAR-100 Aircraft Food-101 Mean Acc. (%) Mean Params. (M)
Full fine-tuning 89.12 70.93 90.96 83.67 85.80
Linear probing 85.95 45.06 88.14 73.05 0.00
BitFit [80] 91.69 68.71 89.59 83.33 0.10
AdaptFormer [7] 91.86 71.71 90.89 84.82 1.19
SparseAdapter [30] 91.20 67.15 89.37 82.57 0.30
MoSA (Ours) 92.22 72.14 91.17 85.18 1.19
@@
General large-scale classification performance. To further evaluate the generality of our method, we compare MoSA with other fine-tuning methods on 3 general classification tasks, as shown in Table 3. MoSA outperforms all baselines, including full fine-tuning, on all datasets. Specifically, compared to AdaptFormer and full fine-tuning, MoSA achieves an average accuracy improvement of 1.51% and 0.36%, respectively, without introducing any additional parameters. It is worth noting that SparseAdapter performs poorly on this benchmark, exhibiting an accuracy drop of 2.25% compared to the standard adapter. This is attributed to the fact that the three datasets in GICD contain relatively sufficient training data, mitigating the advantages of sparsity. However, our method, through multiple sparse adapter experts, demonstrates robust performance in scenarios with both abundant and limited training data, outperforming SparseAdapter by 2.61%, which proves the soundness of our design.
4.3 Extended Results on Different Backbone Scales and LoRA
In this section, we validate the performance of MoSA on different backbone scales and PEFT methods. More experiments on different model architectures (e.g. Swin Transformer) and adapter structures can be found in Appendix.
Table 4: Results on FGVC with ViT-L/16 backbone pre-trained on ImageNet-21K
Full fine-tuning 88.3 85.9 96.7 93.1 86.8 90.16 303.30
Linear probing 84.7 78.9 97.4 89.6 55.1 81.14 0.00
BitFit [80] 88.5 86.4 98.8 93.5 83.2 90.08 0.27
AdaptFormer-64 [7] 89.3 86.3 98.8 93.2 80.9 89.70 3.17
SparseAdapter-64 [30] 89.4 86.8 99.0 93.9 82.3 90.28 0.79
MoSA-64 (Ours) 89.7 87.2 99.4 94.5 84.9 91.06 3.17
MoSA on different backbone scales. Here we evaluate MoSA with a larger backbone ViT-L/16 (303.3M vs. 85.8M ViT-B/16). The results in Table 4 indicate that, with the increase in backbone scale, the performance of the standard adapter cannot surpass full fine-tuning (89.70% vs. 90.16%). Sparse training could improve the performance of adapters , leading by 0.58% and 0.12% over the standard adapter and full fine-tuning, respectively. Our MoSA further enhances the performance by 0.78%, outperforming full fine-tuning by a large margin. This experiment thoroughly demonstrates the importance of sparse training as the scale of pre-trained models increases.
Table 5: Results on FGVC with LoRA
LoRA-16 [34] 87.2 83.5 98.6 89.3 83.7 88.44 0.59
SparseLoRA-16 [30] 87.4 84.9 98.9 91.1 79.9 88.44 0.15
MoSL-16 (Ours) 89.0 85.6 99.3 91.8 83.9 89.92 0.59
MoSA on different PEFT methods. To demonstrate the generality of our approach, we choose another widely used PEFT method, namely LoRA [34]. LoRA achieves parameter-efficient fine-tuning by applying a low-rank decomposition to the weight updates of linear layers. During inference, the additional modules could be merged into the pre-trained model parameters, resulting in zero extra overhead during inference. In Table 5, we compare the LoRA version sparse adapter (namely SparseLoRA) and our MoSA (namely MoSL) with the standard LoRA, with the rank of all LoRA modules set to 16. Our MoSL outperforms both the standard and sparse versions of LoRA, achieving an improvement of 1.48%, further confirming the effectiveness and rationality of our design.
@@ 4.4 Ablation Study
We conduct comprehensive ablation studies to verify the effectiveness of each component in our MoSA design. All ablation experiments are performed on FGVC with ViT-B backbone, and the performances of different strategies are measured using mean accuracy over the datasets. The best component choice, which is also used in the main results, is highlighted with a green background.
Table 6: Hierarchical sparse strategy
Table 7: Consistency regularization
@@
Table 7: Consistency regularization
Hierarchical sparse strategy. The MoE system increases the model capacity, while the sparse gating mechanism also leads to a decrease in the amount of data each expert encounters, resulting in suboptimal performance particularly when training data for downstream tasks is limited. In order to address this issue, we propose a hierarchical sparse strategy, and the results in Table 7 demonstrate the effectiveness of this design. Applying the sparse strategy to both the down-projection and up-projection layer only achieves an accuracy of 88.91%, slightly (0.33%) ahead of Adaptformer (dense down-projection and up-projection layer). However, with the hierarchical sparse strategy, we preserve the down-projection layer as a dense matrix while sparsely splitting the up-projection layer. In this way, MoSA outperforms Adaptformer by a large margin (89.90% vs. 88.58%), demonstrating the importance of this hierarchical strategy.
Table 8: Different alignment strategies
Table 9: Impact of expert number
Table 10: Different inference mechanisms and corresponding efficiency
Alignment position none shallow deep all
Mean Acc. (%) 89.25 89.90 88.84 88.79
@@
Mechanism FLOPs Mean Acc. (%)
fixed 1×1$\times$1 × 88.42
stochastic 1×1$\times$1 × 89.27
ensemble $N\times$ 88.63
merge 1×1$\times$1 × 89.90
Table 9: Impact of expert number
Table 10: Different inference mechanisms and corresponding efficiency
Consistency regularization and feature alignment. The mechanism of stochastic activation may lead to significant discrepancies among different experts, which can have adverse effects on the sparse module merging. In Table 7, we investigate the impact of regularization constraints in the training of MoSA. Without any regularization, our method achieves an accuracy of 88.93%. Applying consistency regularization on the final outputs of the model results in an improvement of 0.32%. And deep alignment of features extracted by different experts can further improve the accuracy by 0.65%. We also explore the impact of different feature alignment positions on model performance. As shown in Table 10, performing alignment solely at shallow layers (closer to the input) can lead to a 0.65% improvement. In contrast, executing alignment at deep layers (closer to the output) results in a 0.41% decrease, attributable to sub-module collapse.
Expert number. To evaluate the compatibility between sparse training and MoE, we vary the number of splits in the adapters of MoSA during training, ranging from 1 to 5 and 8. The results are presented in Table 10. When the number of splits is set to 1, our method degrades to AdaptFormer. We can see that when the number of experts is between 2 and 5, indicating a sparsity level between 20% and 50% for the adapters, MoSA consistently achieves good performance. However, when the number of experts increases to 8, the performance experiences a noticeable decline due to excessive sparsity.
Merging vs. ensembling. During inference, we merge multiple sparse adapters to form a complete one. However, it has been pointed out that stochastic activation during inference can also achieve good performance [86]. So we compare various inference methods, including fixed activation, stochastic activation, logits ensembling, and parameter merging. Results are presented in Table 10. It can be observed that fixed activation has the lowest accuracy, reaching only 88.41% on average. In comparison, stochastic activation achieves an improvement of 0.85%. Logits ensembling also leads to a certain improvement over fixed activation, but the increase of only 0.21% comes at the cost of $N$ times computational complexity ($N$ refers to the expert number), significantly reducing the inference speed. Finally, our parameter merging method achieves the best performance without any additional computation, further enhancing 0.63% over stochastic activation.
Figure 3: Impact of adapter bottleneck dimensions
Figure 3: Impact of adapter bottleneck dimensions
Figure 4: t-SNE visualization on CIFAR-100
Adapter bottleneck dimension. Figure 4 shows the impact of bottleneck dimensions of the adapter in AdaptFormer and MoSA. Overall, the performance shows an increasing trend followed by a decline as the number of trainable parameters increases. Across all bottleneck dimensions, MoSA consistently outperforms AdaptFormer, and our method can even surpass full fine-tuning with a bottleneck dimension of only 16. It is worth noting that the performance of AdaptFormer starts to decline when the bottleneck dimension exceeds 64, while in our method, this turning point occurs at 128. This indicates that our design can more effectively utilize all parameters within the adapter.